# Automated Modeling Method for Pathloss Model Discovery

### v3

@@ Abstract
Modeling propagation is the cornerstone for designing and optimizing next-generation wireless systems, with a particular emphasis on 5G and beyond technologies, such as reconfigurable intelligent surfaces, terahertz (THz) communications, and beam steering in massive multi-input multi-output (mMIMO) systems. Traditional modeling methods have long relied on statistic-based techniques to characterize propagation behavior across different environments. With the expansion of wireless communication systems, there is a growing demand for methods that guarantee the accuracy and interpretability of modeling. Artificial intelligence (AI)-based techniques, in particular, are increasingly being adopted to overcome this challenge. Inspired by recent advancements in AI, this paper proposes a novel approach that accelerates the discovery of path loss models while maintaining interpretability. The proposed method automates the formulation, evaluation, and refinement of the model, facilitating the discovery of the model. We examine two techniques: one based on Deep Symbolic Regression, offering full interpretability, and the second based on Kolmogorov-Arnold Networks, providing two levels of interpretability. Both approaches are evaluated on two synthetic and two real-world datasets. Our results show that Kolmogorov-Arnold Networks achieve the coefficient of determination value $R^{2}$ close to 1 with minimal prediction error, while Deep Symbolic Regression generates compact models with moderate accuracy. Moreover, on the selected examples, we demonstrate that automated methods outperform traditional methods, achieving up to 75% reduction in prediction errors, offering accurate and explainable solutions with potential to increase the efficiency of discovering next-generation path loss models.
@@ I Introduction
Discovering accurate and interpretable scientific formulas that align with established theories is a fundamental goal in science. Traditionally, scientists formulate a hypothesis, derive formulas manually to test it, relying exclusively on theoretical knowledge or statistical characteristics, and subsequently verify its accuracy through simulations and experiments. However, this manual approach, which excludes experimental data during the discovery phase, can be time-consuming. Automated methods have emerged to overcome this limitation by incorporating data directly into the discovery process in several disciplines  [Jumper, Schmidt]. Thus, accurate and interpretable formulas are generated, significantly improving both the speed and quality of the modeling process compared to traditional methods. For instance, recent work in high-energy physics [nour2024] has demonstrated that symbolic regression (SR) can derive models from experimentally measured data that closely resemble the Tsallis distribution, a well-known statistical distribution. Another application lies in improving localization in various domains such as wireless communications, robotics [wang2025], and the Internet of Things (IoT) [shahbazian].
@@
Unlike AI scientists [lu2024ai], that aim to even replace humans in the scientific discovery process, automated methods aim to support the human user to be able to discover interpretable relationships directly from data, without relying on predefined models, by seeking functional representations that capture the underlying patterns [makke2024interpretable]. By leveraging tools such as evolutionary strategies [koza1994] and neural network architectures [Petersen], these methods can efficiently model complex relationships.
@@
Understanding specific electromagnetic propagation models is paramount for wireless networks that have been engineered over time by humans and enable our everyday communication needs. Our ability to build and understand these networks hinges on understanding and modeling how signals are attenuate when propagating through space in realistic environments [Phillips]. Pathloss (PL) models describe the attenuation behavior of signal propagation in various wireless communication systems, incorporating essential factors such as frequency, distance, and presence of various obstacles along the path [Sun]. Over the years, a variety of methods have been proposed for modeling PL, including empirical, deterministic, and machine learning-based approaches, each with its strengths and limitations [Phillips, Ethier]. In recent years, such models have been modeled with deep neural networks (DNNs) [Zeng], in some cases showing better results than some existing analytical models and classical machine learning-based models [Ethier]. However, models based on DNNs are inherently difficult to interpret and explain [dwivedi2023explainable], which hinders the ability to teach humans the principles and intuitions behind models.
In order to enable scientists and engineers to have more efficient PL model discovery tools while maintaining human interpretability, in this paper, we propose a novel approach that accelerates the discovery of PL models (AutoPL modeling method). Figure 1 illustrates the flow diagram of the AutoPL modeling method. Unlike traditional manual methods, where experts formulate models based on established theories and iteratively evaluate and refine them through hands-on analysis, AutoPL method leverages data to formulate the models automatically. These models are then evaluated and refined in an automated manner. The dashed line in the diagram denotes limited iteration.
Motivated by the observation that PL models often involve logarithmic functions in their expression [Phillips], and recognizing that such functions can present challenges for certain model discovery techniques [makke2024interpretable], we investigate an alternative method with different levels of interpretability. The contributions of this paper are outlined as follows.
@@
    We propose a new PL model discovery approach (AutoPL) that automates the model formulation, evaluation, and refinement, facilitating model discovery. In particular, we consider Deep Symbolic Regression (DSR), offering full interpretability, and Kolmogorov-Arnold Networks (KANs), providing two flavors of interpretability, as candidate automation techniques.
@@
    On the selected examples, we demonstrate that automated methods outperform traditional methods, achieving up to 75% reduction in prediction errors, offering accurate and explainable solutions with the potential to increase the efficiency of discovering path loss models for next-generation wireless communication systems.
@@ II-A Automating Scientific Discovery
Perhaps the most recent trend in automating scientific discovery is to develop AI systems that can automate every step of the process, including idea generation, full experiment iteration, and paper writing [lu2024ai]. Slightly more developed, but still in the very early stages of research, various approaches to automating hypothesis generation and verifications have been proposed [cory2024evolving].
@@
SR is one of the earliest and most widely used modeling techniques. A breakthrough came with the introduction of genetic programming (GP), which refines mathematical expressions using evolutionary operations such as mutation, crossover, and selection[koza1994]. GP has been extensively used in SR and remains a dominant approach due to its ability to discover compact expressions [Fortin]. However, GP has significant computational challenges, particularly when searching through large, high-dimensional function spaces [Kronberger]. As the complexity of mathematical expressions increases, the search space expands exponentially, resulting in challenges such as local optima, slow convergence, and insufficient global search guidance. These difficulties have driven the development of hybrid approaches that integrate deep learning and reinforcement learning (RL) with SR, with the aim of improving efficiency and exploration.
@@
Mundhenk et al.[Mundhenk] introduced a neural-guided genetic programming (NGGP) approach, where a recurrent neural network generates initial populations for GP. Unlike standard GP, NGGP uses a trained neural network to initialize the GP search with higher-quality candidate expressions, leading to faster convergence. A closely related approach, Genetic Expert-Guided Learning (GEGL) [Ahn], combines priority queues and GP to select high-quality expressions for RL. However, GEGL uses only one evolutionary step per training iteration, which limits the depth of genetic refinement.
@@
Other hybrid models integrate RL to adjust mutation probabilities, guide selection criteria, or dynamically optimize hyperparameters [Igel, Topchy]. These approaches introduce mechanisms to bias evolutionary search toward higher-reward expressions, helping GP escape local optima while maintaining structural consistency in generated formulas. Another class of hybrid approaches interchanges populations between GP and neural models. Pourchot et al. [Pourchot] and Khadka et al. [Khadka] applied population interchange in deep reinforcement learning tasks such as robotic control. These approaches demonstrate that GP-generated expressions can be used to refine neural network policies, while neural models can improve the diversity of GP populations.
@@
Recent advances in deep learning have introduced alternative methods for SR. Kusner et al. [Kusner] proposed GrammarVAE, which uses a variational autoencoder to model symbolic expressions as latent-space representations while enforcing syntactic constraints. However, the method struggles with recovering exact expressions and often generates invalid mathematical structures. The approach proposed in [Sahoo] incorporates SR into a neural network architecture, where activation functions represent symbolic operators. This approach requires strong assumptions about the functional form of the target equation, which limit its flexibility. A major advancement came with AI Feynman [Udrescu], which combines neural networks with symbolic rule discovery to simplify expressions before applying SR. AI Feynman successfully identifies symmetries and separable functions in data, allowing it to break down complex problems into smaller, easier-to-learn subproblems. However, it does not perform a direct symbolic search, making it unsuitable for some real-world applications.
@@
RL-based approaches have shown notable improvements, particularly in terms of accuracy and convergence speed. One of the key challenges in RL is the efficient sampling of promising expressions. Importance sampling, first introduced by Glynn & Iglehart [Glynn], has been adapted in RL-based approaches to prioritize high-reward expressions while reducing computational cost. Another powerful optimization technique is the cross-entropy method (CEM) [DeBoer], which has been widely used in RL for sampling high-quality solutions from a probabilistic distribution. These approaches have helped address the exploration-exploitation trade-off, ensuring meaningful expressions without excessive random exploration.
@@
KANs [Liukan] represent a significant advancement and shift in neural network design by enhancing the ability to model complex data relationships while maintaining high accuracy and interpretability. The KANs architecture plays a crucial role in achieving these benefits across various domains. Xu et al. [Xu] applied KANs to time series forecasting, introducing T-KAN and MT-KAN, which demonstrated improved predictive performance by effectively capturing temporal dependencies. De Carlo et al[DeCarlo] developed the Graph Kolmogorov-Arnold Network (GKAN), integrating spline-based activation functions into graph neural networks, leading to enhanced accuracy and interpretability in tasks like node classification and link prediction. Bozorgasl & Chen [Bozorgasl] proposed Wav-KAN, incorporating wavelet functions into KANs to efficiently capture both high and low frequency components, improving accuracy and robustness, particularly for multiresolution analysis. By design, KANs enable more interpretability compared to other DNNs as they learn the composition of splines that approximate the data. By looking at the learned graph and the splines at each node, one can interpret what has been learned from the data and why. More recently, they have been extended with symbolic functionality where an actual expression can be extracted from the graph by approximating the learned splines with known functions [liu2024kan].
II-B AI for Pathloss
@@
To date, our preliminary PL approximation study is the first work to consider interpretable AI modeling of wireless networks [anaqreh]. Several works on explainable AI for various aspects of 6G have been proposed, as surveyed in [brik2024explainable]. In some cases, the developed models also claim interpretability, or ”local” interpretability, however, depending on the work, the term refers either to a visual illustration of how much importance a model assigns to certain features or how a deep learning network activates, i.e., activation maps. Of course, certain classical machine learning models, also referred to as ”transparent models” in [brik2024explainable], such as regression trees or linear regression, are also interpretable; however, they tend to yield less accurate models.
@@
With respect to AI-aided PL modeling, [Zeng] proposed a meta-learning approach based on DNNs to develop a PL prediction model when only sparse measurements are available. However, its superiority compared to conventional techniques is not immediately clear. On the other hand, [Ethier] showed that a fully connected network was superior to a boosted tree model and two conventional approaches in several scenarios.
@@ III Problem statement
The guiding hypothesis of this work is the possibility of automating the formulation, evaluation, and refinement steps of the conventional PL model development process. By doing so, the aim is to equip scientists and engineers with more efficient tools for discovering PL models and, at the same time, maintaining human interpretability. In order to validate this hypothesis, we consider two classes of models: analytical and empirical. The two considered analytical models are $PL^{ABG}$ (Alpha-Beta-Gamma) and $PL^{CI}$ (Close-in) models [Sun2], which are widely used in modeling PL in wireless systems. The two empirical models, herein referred to as $PL^{EI}$ and $PL^{EO}$, correspond to indoor and outdoor environments, respectively, and have been developed from measurements in  [ElChall]. As analytical baselines for corresponding PL models, we include the indoor multiwall-and-floor ($PL^{MWF}$) and outdoor free space ($PL^{FS}$) models. Based on the data generated with the analytical and baseline models, as well as the empirically measured one, we aim to automatically learn the most likely expression, i.e., formulas, that best describe the respective PL.
@@ III-A ABG and CI Analytical Models
where $PL^{\text{ABG}}(f,d)$ is the path loss in dB, $\alpha$ and $\gamma$ are coefficients that represent the dependence on distance ($d$) and frequency ($f$) respectively. $\beta$ is an optimized offset value for path loss, and $\chi_{\sigma}^{\text{ABG}}$ is a random variable that represents shadow fading, which takes a random value following the Gaussian distribution with zero mean and a standard deviation of $\sigma$.
The equation describes the CI model:
@@
where $n$ is the path loss exponent (PLE), $d$ is the transmitter-receiver distance in three dimensions, $FSPL(f,1\,\text{m})$ is the loss of the path in free space at 1 meter and $\chi_{\sigma}^{\text{CI}}$ represents the shadow fading, described as a random variable following a Gaussian distribution with a mean of zero and a standard deviation of $\sigma$.

### v4

@@ Abstract
Modeling propagation is the cornerstone for designing and optimizing next-generation wireless systems, with a particular emphasis on 5G and beyond technologies, such as reconfigurable intelligent surfaces, terahertz (THz) communications, and beam steering in massive multi-input multi-output (mMIMO) systems. Traditional modeling methods have long relied on statistic-based techniques to characterize propagation behavior across different environments. With the expansion of wireless communication systems, there is a growing demand for methods that guarantee the accuracy and explicit functional interpretability of modeling. Artificial intelligence (AI)-based techniques, in particular, are increasingly being adopted to overcome this challenge. Inspired by recent advancements in AI, this paper proposes a novel approach that accelerates the discovery of pathloss (PL) models while maintaining interpretability, referring to explicit functional interpretability in the remainder of the paper unless otherwise stated. The proposed method automates the formulation, evaluation, and refinement of the model, facilitating the discovery of the model. We examine two techniques: one based on Deep Symbolic Regression (DSR), offering interpretability, and the second based on Kolmogorov-Arnold Networks (KANs), providing two levels of interpretability. Both approaches are evaluated on two synthetic and two real-world datasets. Our results show that KANs achieve the coefficient of determination value $R^{2}$ close to 1 with minimal prediction error, while DSR generates compact models with moderate accuracy. Moreover, on the selected examples, we demonstrate that automated methods outperform traditional methods, achieving up to 75% reduction in prediction errors, offering accurate and functionally interpretable and analyzable solutions with potential to increase the efficiency of discovering next-generation pathloss models.
@@ I Introduction
Emerging AI for Scientific Discovery: Discovering accurate and interpretable scientific formulas that align with established theories is a fundamental goal in science. Traditionally, scientists formulate a hypothesis, derive formulas manually to test it, relying exclusively on theoretical knowledge or statistical characteristics, and subsequently verify its accuracy through simulations and experiments. However, this manual approach, which excludes experimental data during the discovery phase, can be time-consuming. Automated methods have emerged to overcome this limitation by incorporating data directly into the discovery process in several disciplines  [23, 39]. Thus, accurate and interpretable formulas are generated, significantly improving both the speed and quality of the modeling process compared to traditional methods. For instance, recent work in high-energy physics [32] has demonstrated that symbolic regression (SR) can derive models from experimentally measured data that closely resemble the Tsallis distribution, a well-known statistical distribution. Another application lies in improving localization in various domains such as wireless communications, robotics [45], and the Internet of Things [40].
@@
Unlike AI scientists [31], who aim to even replace humans in the scientific discovery process, automated methods aim to support the human user to be able to discover interpretable relationships directly from data, without relying on predefined models, by seeking functional representations that capture the underlying patterns [33]. By leveraging tools such as evolutionary strategies [25] and neural network architectures [35], these methods can efficiently model complex relationships.
@@
AI for PL Modeling: Understanding specific electromagnetic propagation models is paramount for wireless networks that have been engineered over time by humans and enable our everyday communication needs. Our ability to build and understand these networks hinges on understanding and modeling how signals are attenuated when propagating through space in realistic environments [36]. Pathloss (PL) models describe the attenuation behavior of signal propagation in various wireless communication systems, incorporating essential factors such as frequency, distance, and the presence of various obstacles along the path [42]. Over the years, a variety of methods have been proposed for modeling PL, including empirical, deterministic, and machine learning-based approaches, each with its strengths and limitations [36, 17]. The traditional empirical models, such as the widely used Alpha-Beta-Gamma (ABG) and Close-In (CI) models, rely on rigid, predefined forms. This structural inflexibility limits their ability to capture the complex environmental effects, including the dual-slope attenuation characteristic of urban streets [13] or the molecular absorption peaks inherent to Terahertz (THz) communications [21].
Need for Functional Interpretability: In recent years, such models have been modeled with deep neural networks (DNNs) [48], in some cases showing better results than some existing analytical models and classical machine learning-based models [17]. However, models based on DNNs are inherently difficult to interpret and explain [16], which hinders the ability to teach humans the principles and intuitions behind models. Unlike the traditional, physics-based methods, DNNs do not provide an explicit mathematical expression to describe the relationship between parameters (e.g., frequency, distance) and the signal attenuation. This lack of interpretability, referring to explicit functional interpretability in the remainder of the paper unless otherwise stated, makes it difficult for future engineers or researchers to validate the model against physical laws or derive generalizable insights for network planning, essentially trading physical understanding for numerical accuracy.
To bridge the aforementioned gap, we propose AutoPL, a novel approach that accelerates the discovery of PL models. Unlike traditional methods where experts manually formulate and refine models based on established theories, AutoPL automates the symbolic discovery process to retrieve interpretable mathematical expressions directly from data (see Figure 1). Motivated by the observation that PL models often involve logarithmic functions [36], which can present challenges for certain discovery techniques [33], we investigate this data-driven alternative to ensure explicit functional interpretability. In this context, interpretability refers to the availability of explicit mathematical expressions and transparent parameter dependencies, rather than a direct interpretation of underlying electromagnetic propagation mechanisms. The contributions of this paper are outlined as follows.
@@
    We propose a new PL model discovery approach (AutoPL) that automates the model formulation, evaluation, and refinement, facilitating model discovery. In particular, we consider Deep Symbolic Regression (DSR), offering explicit functional interpretability, and Kolmogorov-Arnold Networks (KANs), providing two flavors of explicit functional interpretability, as candidate automation techniques.
@@
    On the selected examples, we demonstrate that automated methods outperform traditional methods, achieving up to 75% reduction in prediction errors, offering accurate and explainable solutions with the potential to increase the efficiency of discovering PL models for next-generation wireless communication systems.
@@ II-A Automating Scientific Discovery
Perhaps the most recent trend in automating scientific discovery is to develop AI systems that can automate every step of the process, including idea generation, full experiment iteration, and paper writing [31]. Slightly more developed, but still in the very early stages of research, various approaches to automating hypothesis generation and verifications have been proposed [12].
@@
SR is one of the earliest and most widely used modeling techniques. A breakthrough came with the introduction of genetic programming (GP), which refines mathematical expressions using evolutionary operations such as mutation, crossover, and selection[25]. GP has been extensively used in SR and remains a dominant approach due to its ability to discover compact expressions [18]. However, GP has significant computational challenges, particularly when searching through large, high-dimensional function spaces [26]. As the complexity of mathematical expressions increases, the search space expands exponentially, resulting in challenges such as local optima, slow convergence, and insufficient global search guidance. These difficulties have driven the development of hybrid approaches that integrate deep learning (DL) and reinforcement learning (RL) with SR, with the aim of improving efficiency and exploration.
@@
Mundhenk et al.[34] introduced a neural-guided genetic programming (NGGP) approach, where a recurrent neural network generates initial populations for GP. Unlike standard GP, NGGP uses a trained neural network to initialize the GP search with higher-quality candidate expressions, leading to faster convergence. A closely related approach, Genetic Expert-Guided Learning (GEGL) [2], combines priority queues and GP to select high-quality expressions for RL. However, GEGL uses only one evolutionary step per training iteration, which limits the depth of genetic refinement.
@@
Other hybrid models integrate RL to adjust mutation probabilities, guide selection criteria, or dynamically optimize hyperparameters [22, 43]. These approaches introduce mechanisms to bias evolutionary search toward higher-reward expressions, helping GP escape local optima while maintaining structural consistency in generated formulas. Another class of hybrid approaches interchanges populations between GP and neural models. Pourchot et al. [37] and Khadka et al. [24] applied population interchange in deep reinforcement learning tasks such as robotic control. These approaches demonstrate that GP-generated expressions can be used to refine neural network policies, while neural models can improve the diversity of GP populations.
@@
Recent advances in DL have introduced alternative methods for SR. Kusner et al. [27] proposed GrammarVAE, which uses a variational autoencoder to model symbolic expressions as latent-space representations while enforcing syntactic constraints. However, the method struggles with recovering exact expressions and often generates invalid mathematical structures. The approach proposed in [38] incorporates SR into a neural network architecture, where activation functions represent symbolic operators. This approach requires strong assumptions about the functional form of the target equation, which limit its flexibility. A major advancement came with AI Feynman [44], which combines neural networks with symbolic rule discovery to simplify expressions before applying SR. AI Feynman successfully identifies symmetries and separable functions in data, allowing it to break down complex problems into smaller, easier-to-learn subproblems. However, it does not perform a direct symbolic search, making it unsuitable for some real-world applications.
@@
RL-based approaches have shown notable improvements, particularly in terms of accuracy and convergence speed. One of the key challenges in RL is the efficient sampling of promising expressions. Importance sampling, first introduced by Glynn & Iglehart [19], has been adapted in RL-based approaches to prioritize high-reward expressions while reducing computational cost. Another powerful optimization technique is the cross-entropy method (CEM) [5], which has been widely used in RL for sampling high-quality solutions from a probabilistic distribution. These approaches have helped address the exploration-exploitation trade-off, ensuring meaningful expressions without excessive random exploration.
@@
KANs [30] represent a significant advancement and shift in neural network design by enhancing the ability to model complex data relationships while maintaining high accuracy and explicit functional interpretability. The KANs architecture plays a crucial role in achieving these benefits across various domains. Xu et al. [47] applied KANs to time series forecasting, introducing T-KAN and MT-KAN, which demonstrated improved predictive performance by effectively capturing temporal dependencies. De Carlo et al[14] developed the Graph Kolmogorov-Arnold Network (GKAN), integrating spline-based activation functions into graph neural networks, leading to enhanced accuracy and interpretability in tasks like node classification and link prediction. Bozorgasl & Chen [6] proposed Wav-KAN, incorporating wavelet functions into KANs to efficiently capture both high and low frequency components, improving accuracy and robustness, particularly for multiresolution analysis. By design, KANs enable more interpretability compared to other DNNs as they learn the composition of splines that approximate the data. By looking at the learned graph and the splines at each node, one can interpret what has been learned from the data and why. More recently, they have been extended with symbolic functionality where an actual expression can be extracted from the graph by approximating the learned splines with known functions [29].
II-B AI for PL
@@
To date, our preliminary PL approximation study [3] published in the same month with [15] are the first work to consider interpretable AI modeling of wireless networks. While [3] focused on KANs and DSR with the ABG and CI models, [15] focused on symbolic regression through an evolutionary algorithm and ITU-R P.1546-5. As this paper was available as pre-print and underwent revision, KAN-based modeling has also been considered for radio map prediction [28]. Several works on explainable AI for various aspects of 6G have been proposed, as surveyed in [7]. In some cases, the developed models also claim interpretability, or ”local” interpretability, however, depending on the work, the term refers either to a visual illustration of how much importance a model assigns to certain features or how a DL network activates, i.e., activation maps. Of course, certain classical machine learning models, also referred to as ”transparent models” in [7], such as regression trees or linear regression, are also interpretable; however, they tend to yield less accurate models.
@@
With respect to AI-aided PL modeling, [48] proposed a meta-learning approach based on DNNs to develop a PL prediction model when only sparse measurements are available. However, its superiority compared to conventional techniques is not immediately clear. On the other hand, [17] showed that a fully connected network was superior to a boosted tree model and two conventional approaches in several scenarios.
@@ III Problem statement
The guiding hypothesis of this work is the possibility of automating the formulation, evaluation, and refinement steps of the conventional PL model development process. By doing so, the aim is to equip scientists and engineers with more efficient tools for discovering PL models and, at the same time, maintaining human interpretability through mathematical expressions. In order to validate this hypothesis, we consider two classes of models: analytical and empirical. The two considered analytical models are $PL^{ABG}$ (Alpha-Beta-Gamma) and $PL^{CI}$ (Close-in) models [41], which are widely used in modeling PL in wireless systems. The two empirical models, herein referred to as $PL^{EI}$ and $PL^{EO}$, correspond to indoor and outdoor environments, respectively, and have been developed from measurements in  [8]. As analytical baselines for corresponding PL models, we include the indoor multiwall-and-floor ($PL^{MWF}$) and outdoor free space ($PL^{FS}$) models. Based on the data generated with the analytical and baseline models, as well as the empirically measured one, we aim to automatically learn the most likely expression, i.e., formulas, that best describe the respective PL.
Similar to established PL models such as CI and ABG, which abstract complex electromagnetic interactions into compact empirical laws, the flow proposed in Figure 1 aims to speed up the development of models that prioritize analytical usability and generalization over microscopic physical decomposition.
@@ III-A ABG and CI Analytical Models
where $PL^{\text{ABG}}(f,d)$ is the PL in dB, $\alpha$ and $\gamma$ are coefficients that represent the dependence on distance ($d$) and frequency ($f$) respectively. $\beta$ is an optimized offset value for PL, and $\chi_{\sigma}^{\text{ABG}}$ is a random variable that represents shadow fading, which takes a random value following the Gaussian distribution with zero mean and a standard deviation of $\sigma$.
The following equation describes the CI model:
@@
where $f$ is the frequency in Hz, $n$ is the PL exponent (PLE), $d$ is the transmitter-receiver distance in three dimensions, $FSPL(f,1\,\text{m})$ is the loss of the path in free space at 1 meter and $\chi_{\sigma}^{\text{CI}}$ represents the shadow fading, described as a random variable following a Gaussian distribution with a mean of zero and a standard deviation of $\sigma$.